# Communication Cost Reduction for Subgraph Counting under Local Differential Privacy via Hash Functions

### v1

@@ II-D Smooth Sensitivity
Simply put, the local sensitivity of $f$ on $\{0,1\}^{n}$ measures the largest change in $f$ due to a single edge alteration when starting with the adjacency list $a$. The sensitivity of $f$ at distance $k$ expands on this idea by looking at the changes resulting from a single alteration in the adjacency lists that are within $k$ modifications from $a$.
Definition 6 (Sensitivity at distance $k$ from Definition 3.1 of [40]).
@@
Definition 7 (Smooth Sensitivity from Definition 2.2 of [40]).
@@
Definition 8 (Smooth Sensitivity mechanism from Lemma 2.5 of [40]).
@@
This result was later improved in [41] who proved that $\mathcal{R}(a)$ could be changed to $f(a)+\frac{2(\gamma-1)}{\varepsilon}S^{*}_{f,\varepsilon/2(\gamma-1)}(a)\cdot Z_{\gamma}$.
@@
III Previous Works on Subgraph Countings under Edge Local Differential Privacy
@@
Let us suppose the the original graph is $G=(V,E)$ when $V=\{v_{1},\dots,v_{n}\}$ is a set of nodes or users and $E\subseteq\{\{u,v\}:u,v\in V\}$ is a set of edges or relationships. We can publish $G$ using the randomized response mechanism defined in the previous section. Given that $a^{\prime}_{i,j}=RR(a_{i,j})$ obtained from the randomized response, we can define a graph $G^{\prime}=(V,E^{\prime})$ with its adjacency matrix given by $A^{\prime}=(a^{\prime}_{i,j})_{1\leq i,j\leq n}$. While one can release the count of subgraphs by counting them in $G^{\prime}$, such a release is known to introduce a significant error [22].
III-A Two-Step Mechanism
@@
To mitigate the error, Imola, Murakami, and Chaudhuri [20] introduced a two-step mechanism to release the count of triangles. A tuple of three nodes, represented as $(v_{i},v_{j},v_{k})$ where $i>j>k$ is a triangle if $\{v_{i},v_{j}\},\{v_{i},v_{k}\},\{v_{j},v_{k}\}\in E$. We denote the set of triangles in $G$ by $\Delta_{G}$. Furthermore, for tuples in $\Delta_{G}$ where the initial element is $v_{i}$, we denote them as $\Delta_{i}$. We have that $|\Delta_{G}|=\sum\limits_{i}|\Delta_{i}|$.
@@
The first step, represented by $\mathcal{M}_{1}$, involves releasing the obscured graph $A^{\prime}$. In the second step, represented by $\mathcal{M}_{2}$, every user $v_{i}$ provides an estimated value for $|\Delta_{i}|$ that is based on both the real adjacency vector $a_{i}$ and the distorted graph $G^{\prime}$. Let $\Delta^{\prime}_{i}=\{(v_{j},v_{k}):\{v_{i},v_{j}\},\{v_{i},v_{k}\}\in G,\{v_{j},v_{k}\}\in G^{\prime},\text{ and }i>j>k\}$. Consider a function $f$ defined as $f(a_{i})=|\Delta^{\prime}_{i}|$. The randomized funtion of $\mathcal{M}_{2}$, represented as $\mathcal{R}_{2}$, aligns with what is used in the edge local Laplacian mechanism. Specifically, users are asked to add a Laplacian noise into $f(a_{i})$. The mechanism’s aggregator then sums the outcomes derived from these randomized functions.
@@
The two-step mechanism can significantly increase the precision of the triangle counting. However, it has a large variance. Consider $i>i^{\prime}>j>k$ such that $\{v_{i},v_{j}\},\{v_{i},v_{k}\},\{v_{i^{\prime}},v_{j}\},\{v_{i^{\prime}},v_{k}\}\in E$ and $\{v_{j},v_{k}\}\notin E$. If, by the randomized response, $\{v_{j},v_{k}\}\in E^{\prime}$, that would contribute two to the number of triangles. When the input graph has a large number of $C_{4}$, the variance becomes large. The authors of [22] propose a technique called four-cycle trick to reduce the variance. Define $\Delta_{i}^{\prime\prime}:=\{(v_{j},v_{k})\in\Delta_{i}^{\prime}:\{v_{i},v_{k}\}\in G^{\prime}\}$. Instead of having $f(a_{i})=|\Delta_{i}^{\prime}|$, they use $f(a_{i})=|\Delta_{i}^{\prime\prime}|$.
III-B Communication Cost
@@
In many real-world social networks, each user is typically connected to a constant number of other users [42]. This means the number of connections (or ‘ones’ in the adjacency list $a_{i}$) is a lot fewer than $n$. So, instead of sending their entire adjacency list $a_{i}$ to the aggregator, user $v_{i}$ can just share the list of those directly connected to them. This approach reduces the communication cost from $n$ bits down to $\Theta(\log n)$.
@@
The four-cycle trick mentioned in the previous section can help reducing the communication cost. As a triangle $(v_{i},v_{j},v_{k})\in\Delta_{i}^{\prime\prime}$ only if $\{v_{i},v_{k}\}\in E^{\prime}$, user $v_{i}$ can calculate $|\Delta_{i}^{\prime\prime}|$ without accessing the information about whether $\{v_{j},v_{k}\}\in E^{\prime}$ or not when $\{v_{i},v_{k}\}\notin E^{\prime}$. Still, we think that the savings from the trick are not substantial enough to make the two-step method viable for big graphs.
IV Group Randomized Response
In this section, we introduce a mechanism designed for publishing graphs in a distributed environment under the edge local differential privacy. We have termed this approach the group randomized response (GRR). The mechanism is a four-step process:
@@ IV Group Randomized Response
    Degree Sharing Step: In this step, each user share their degree protected by the Laplacian mechanism.
@@
    Grouping Step: In this step, nodes and edges are categorized into groups using the linear congruence hashing function.
@@
    Randomized Response Step: Here, obfuscation is applied solely to one randomly chosen representative from each group. Only edges corresponding to the representatives are sent to the aggregator.
  4. 4.
@@
    Counting Step: Subgraphs are counted based on the information from the representatives.
@@
Thanks to the sampling created by the grouping, this mechanism achieves a notably reduced communication cost. Moreover, the privacy budget amplification from the sampling [34] further decreases the probabilities of bit-flipping during the randomized response phase, thus further reduces the cost.
IV-A Degree Sharing Step
@@
During this step, each user $v_{i}$ calculates the count of their connections, specifically |$|\{j<i:\{i,j\}\in E\}|$. They then share this count using the Laplacian mechanism with the privacy budget of $\varepsilon_{0}$. The shared count, denoted by $\tilde{d}_{i}$, will be used only for adjusting bias in the fourth step. If there are a lot of users ($n$ is large), any errors from this count are minor even if a large Laplacian noise is added. Thus, we can choose a relatively small value for $\varepsilon_{0}$.
IV-B Grouping Step
@@
In this step, users do not publish any information. The aim here is to form random groups of edges that are of similar sizes. It is crucial that this randomization can be reproduced, and that the division is easily shareable among all participants. To achieve such a grouping that meets these criteria, we employ the linear congruence hashing, which is discussed in Section II.
[…]
@@
Let $s$ be a parameter called sampling rate, $m$ be an integer such that $m=\lceil p/s\rceil$, and let $h_{i}(j)=h_{\theta_{i},\phi_{i}}(j)=((\theta_{i}\cdot j+\phi_{i})\text{ mod }p)\text{ mod }m$. Each user $v_{i}$ classifies the set of users $\{v_{1},\dots,v_{n}\}$ into $m$ disjoint groups, denoted by $S_{i,1},\dots,S_{i,m}$. The group $S_{i,t}$ can be defined as $\{v_{j}:h_{i}(j)=t\}$. By the property of the linear congruence hashing, we have that $s-1\leq|S_{i,t}|\leq s$. To make sure that all the sets have size $s$, we add a dummy element to all $S_{i,t}$ with size $s-1$. There is no edge between $v_{i}$ and the dummy node.
[…]
@@
We then choose a representative for each group, $S_{i,t}$. Every member of the group has an equal probability of being selected, which is $1/s$. Let us call the chosen member from group $S_{i,t}$ as $v_{c_{t}}$. When $c_{t}$ is chosen from the group containing $k$, we will employ the notation $a_{i,h_{i}(k)}=a_{i,c_{t}}$ for convenience.
IV-C Randomized Response Step
@@
IV-D Counting Step

### v2

@@ 2.3 Basic Mechanisms
The local sensitivity of $f$ on $\{0,1\}^{n}$ measures the largest change in $f$ due to a single edge alteration when starting with the adjacency list $a$. The sensitivity of $f$ at distance $k$ expands on this idea by looking at the changes resulting from a single alteration in the adjacency lists that are within $k$ modifications from $a$.
Definition 6 (Sensitivity at Distance $k$).
@@
Definition 7 (Smooth Sensitivity).
@@
Definition 8 (Smooth Sensitivity Mechanism).
@@
This result was later improved in Yamamoto & Shibuya (2023) who proved that $\mathcal{R}(a)$ could be changed to $f(a)+\frac{2(\gamma-1)}{\varepsilon}S^{*}_{f,\varepsilon/2(\gamma-1)}(a)\cdot Z_{\gamma}$.
@@
3 Previous Mechanism
@@
Let us suppose the the original graph is $G=(V,E)$ when $V=\{v_{1},\dots,v_{n}\}$ is a set of nodes or users and $E\subseteq\{\{u,v\}:u,v\in V\}$ is a set of edges or relationships. We can publish $G$ using the randomized response mechanism defined in the previous section. Given that $a^{\prime}_{i,j}=\textnormal{{RR}}(a_{i,j})$ obtained from the randomized response, we can define a graph $G^{\prime}=(V,E^{\prime})$ with its adjacency matrix given by $A^{\prime}=(a^{\prime}_{i,j})_{1\leq i,j\leq n}$. One can determine the number of subgraphs in $G^{\prime}$ by counting them; however, this method introduces a significant error (Imola et al., 2022a). Specifically, the error magnitude is $\mathcal{O}\left(\frac{C_{4}(G)}{\varepsilon^{2}}+\frac{n^{3}}{\varepsilon^{6}}\right)$, which is large in relation to $n$ and becomes prohibitive as $\varepsilon$ decreases, particularly when aiming for higher levels of privacy.
3.1 Two-Step Mechanism
@@
To mitigate the error, Imola et al. (2021) introduced a two-step mechanism to release the count of triangles. A tuple of three nodes, represented as $(v_{i},v_{j},v_{k})$ where $i>j>k$ is a triangle if $\{v_{i},v_{j}\},\{v_{i},v_{k}\},\{v_{j},v_{k}\}\in E$. We denote the set of triangles in $G$ by $\Delta_{G}$. Furthermore, for tuples in $\Delta_{G}$ where the initial element is $v_{i}$, we denote them as $\Delta_{i}$. We have that $|\Delta_{G}|=\sum\limits_{i}|\Delta_{i}|$.
@@
The first step, represented by $\mathcal{M}_{1}$, involves releasing the obscured graph $G^{\prime}=(V,E^{\prime})$ produced by the randomized response. In the second step, represented by $\mathcal{M}_{2}$, every user $v_{i}$ provides an estimated value for $|\Delta_{i}|$ that is based on both the real adjacency vector $a_{i}$ and the distorted graph $G^{\prime}$. Let $\Delta^{\prime}_{i}=\{(v_{j},v_{k}):\{v_{i},v_{j}\},\{v_{i},v_{k}\}\in E,\{v_{j},v_{k}\}\in E^{\prime},\text{ and }i>j>k\}$. Consider a function $f$ defined as $f(a_{i})=|\Delta^{\prime}_{i}|$. The randomized funtion of $\mathcal{M}_{2}$, represented as $\mathcal{R}_{2}$, aligns with what is used in the edge local Laplacian mechanism. Specifically, users are asked to add a Laplacian noise into $f(a_{i})$. The mechanism’s aggregator then sums the outcomes derived from these randomized functions.
@@
The two-step mechanism can significantly increase the precision of the triangle counting. However, it has a large variance. Consider $i>i^{\prime}>j>k$ such that $\{v_{i},v_{j}\},\{v_{i},v_{k}\},\{v_{i^{\prime}},v_{j}\},\{v_{i^{\prime}},v_{k}\}\in E$ and $\{v_{j},v_{k}\}\notin E$. If, by the randomized response, $\{v_{j},v_{k}\}\in E^{\prime}$, that would contribute two to the number of triangles. When the input graph has a large number of $C_{4}$, the variance becomes large. The authors of Imola et al. (2022a) propose a technique called four-cycle trick to reduce the variance. Define $\Delta_{i}^{\prime\prime}:=\{(v_{j},v_{k})\in\Delta_{i}^{\prime}:\{v_{i},v_{k}\}\in G^{\prime}\}$. Instead of having $f(a_{i})=|\Delta_{i}^{\prime}|$, they use $f(a_{i})=|\Delta_{i}^{\prime\prime}|$. For the rest of the article, we will use ARROne to denote the version of ARR that uses this 4-cycle trick.
3.2 Communication Cost
@@
In many real-world social networks, each user is typically connected to a constant number of other users (Barabási, 2013). This means the number of connections (or ‘ones’ in the adjacency list $a_{i}$) is a lot fewer than $n$. So, instead of sending their entire adjacency list $a_{i}$ to the aggregator, user $v_{i}$ can just share the list of those directly connected to them. This approach reduces the communication cost from $n$ bits down to $\Theta(\log n)$.
@@
The four-cycle trick mentioned in the previous section can help reducing the communication cost. As a triangle $(v_{j},v_{k})\in\Delta_{i}^{\prime\prime}$ only if $\{v_{i},v_{k}\}\in E^{\prime}$, user $v_{i}$ can calculate $|\Delta_{i}^{\prime\prime}|$ without accessing the information about whether $\{v_{j},v_{k}\}\in E^{\prime}$ or not when $\{v_{i},v_{k}\}\notin E^{\prime}$. Still, we think that the savings from the trick are not substantial enough to make the two-step method viable for big graphs.
4 Our Mechanism: Group Randomized Response
In this section, we introduce a mechanism designed for publishing graphs in a distributed environment under the edge local differential privacy. We have termed this approach the group randomized response (GroupRR). The mechanism is a three-step process:
@@ 4 Our Mechanism: Group Randomized Response
    Degree Sharing Step: In this step, each user shares their degree protected by the local Laplacian mechanism.
@@
    Group Randomized Response Step: In this step, a sampled adjacency list is published by each user. First, nodes and edges are categorized into groups using the linear congruence hashing function. Then, one value is sampled from each group and only this value is obfuscated and sent to the central server.
@@
    Counting Step: Subgraphs are counted based on the information from the representatives. Then, each user shares the number of subgraphs protected by the local Laplacian mechanism.
Figure 1 provides an overview of the various steps of the mechanism, while Figure 2 emphasizes the communication between users and the central server.
Figure 2: Details of the interactions between each user and the server during the execution of our mechanism
@@
Thanks to the sampling created by the formation of groups, this mechanism achieves a notably reduced communication cost. Moreover, the privacy budget amplification from sampling (Balle et al., 2018) further decreases the probabilities of bit-flipping during the randomized response phase, further reducing the communication cost.
4.1 Degree Sharing Step
@@
During this step, each user $v_{i}$ calculates the count of their connections, specifically $d_{i}=|\{j<i:\{i,j\}\in E\}|$. We call $d_{i}$ as low degree in this paper. They then share the low degree using the local Laplacian mechanism with the privacy budget of $\varepsilon_{0}$. The shared count, denoted by $\tilde{d}_{i}$, will be used only for adjusting bias in the fourth step. If there are a lot of users ($n$ is large), any errors from this count are minor even if a large Laplacian noise is added. Thus, we can choose a relatively small value for $\varepsilon_{0}$. This step is described as in Algorithm 1.
1
2Function DegreeSharing
 Input: Graph $G=(V,E)$, privacy budget $\varepsilon_{0}$
 Output: Estimated degree $\tilde{d}_{i}$ of each user
3 [User $i$] Calculate and send $\tilde{d}_{i}\leftarrow d_{i}+\textnormal{{Lap}}(\frac{1}{\varepsilon_{0}})$ to the central server
4
Algorithm 1 The first step of our proposed method, degree sharing step
4.2 Group Randomized Response Step
@@
The goal of this step is to publish an obfuscated adjacency list in a communication efficient way. Our solution is to form random groups of edges that are of similar sizes. It is crucial that this randomization can be reproduced, and that the division is easily shareable among all participants. To achieve such a grouping that meets these criteria, we employ the linear congruence hashing, which is discussed in Section 2. Once that grouping is done, one value is sampled per group. This value is then obfuscated with the randomized response mechanism and broadcasted. This step is described in Algorithm 2.
At Line 2 of the algorithm, the central server calculates $p\in\mathbb{N}$, which represents the smallest prime number greater than $n$, the graph’s size. Then, at Line 3, for each node $v_{i}$ when $i\in[|1,n|]$, the server randomly generates two values: $\theta_{i}$ from the range [|1,p1|][|1,p-1|] and $\phi_{i}$ from the range [|0,p1|][|0,p-1|]. These values act as coefficients for the linear congruence hashing. After generating these coefficients, the server shares them, along with $p$, with all users.
@@
Recall the definition of $h_{\mathsf{a},\mathsf{b}}$ in Section 2. Let $s\in\mathbb{N}^{*}$ be a parameter called sampling size which will be the size of each group, $m$ be an integer such that $m=\lceil p/s\rceil$, and let $h_{i}(j)=h_{\theta_{i},\phi_{i}}(j)=((\theta_{i}\cdot j+\phi_{i})\text{ mod }p)\text{ mod }m$. At Line 4, each user $v_{i}$ classifies the set of users $\{v_{1},\dots,v_{n}\}$ into $m$ disjoint groups, denoted by $S_{i,1},\dots,S_{i,m}$. The group $S_{i,t}$ can be defined as $\{v_{j}:h_{i}(j)=t\}$. By the property of the linear congruence hashing, we have that $s-1\leq|S_{i,t}|\leq s$. To make sure that all the sets have size $s$, we add a dummy element to all $S_{i,t}$ with size s1s-1. There is no edge between $v_{i}$ and the dummy node.
@@
At Line 5, we then choose a representative for each group, $S_{i,t}$. Every member of the group has an equal probability of being selected, which is $1/s$. Let us call the chosen member from group $S_{i,t}$ as $v_{c_{t}}$. When $c_{t}$ is chosen from the group containing $k$, we will employ the notation $a_{i,h_{i}(k)}:=a_{i,c_{t}}$ for convenience. Only the value of the representative edge $a_{i,c_{t}}$ is published via the randomized response mechanism, and the actual selected index $c_{t}$ is not shared. We introduce this quantity only to simplify the description and analysis of the algorithm.
It should be emphasized that even though the hashing coefficients are determined by the central server, this does not pose a privacy risk. This is because the hashing functions are employed to form groups that are intentionally public. The privacy-sensitive element is the representative value chosen for each group, and the selection of these representatives does not involve the hash functions.
@@
To ensure that each edge is published only once, each user only publishes connections with users who have a smaller index than their own. To enforce this condition, in Line 6, we assign $a^{\prime}_{i,c_{t}}$ to zero when $c_{t}$ is greater than $i$.
1
2Function GroupRandomizedResponse
 Input: Graph $G=(V,E)$, sampling size $s$, privacy budget $\varepsilon_{1}$
 Output: The grouped and obfuscated adjacency list $(a^{\prime}_{i,c_{1}},\dots,a^{\prime}_{i,c_{m}})$ for each user
3 [Server] Calculate and broadcast $p$ the smallest prime number larger than $n$
4 [Server] For all $i\in[1,n]$, randomly select and broadcast $(\theta_{i},\phi_{i})$ from $1,p-1]\times[0,p-1]$
5 [User $i$] Recall the definition of $h_{\mathsf{a},\mathsf{b}}$ in Section 2. Let $m$ be an integer such that $m=\lceil p/s\rceil$, and let $h_{i}(j)=h_{\theta_{i},\phi_{i}}(j)=((\theta_{i}\cdot j+\phi_{i})\text{ mod }p)\text{ mod }m$. For each $t\in[1,m]$, $S_{i,t}\leftarrow\{v_{j}\mid h_{i}(j)=t\}$;
6 [User $i$] Choose randomly $v_{c_{t}}$ from the set $S_{i,t}$
7 [User $i$] For all $t\in[1,m]$, if $c_{t}>i,a_{i,c_{t}}\leftarrow 0$;
8 [User $i$] Perform the randomized response with privacy budget $\varepsilon^{\prime}=\ln(1+s(e^{\varepsilon_{1}}-1))$ on $(a_{i,c_{1}},\dots,a_{i,c_{m}})$. The result $(a^{\prime}_{i,c_{1}},\dots,a^{\prime}_{i,c_{m}})$ is then broadcasted to all users (We note that only $a^{\prime}_{i,c_{i}}$ for all $1\leq i\leq m$ is published. The index $c_{i}$ is not.);
9
10
Algorithm 2 The second step of our proposed method, group randomized response step
4.3 Counting Step
The final step of our method is outlined in Algorithm 3. In Lines 2-3, we update the value of $a^{\prime}_{i,c_{t}}$ to $\tilde{a}_{i,c_{t}}=\omega_{i}\cdot a^{\prime}_{i,c_{t}}-\tilde{\sigma}_{i}$. In the next section, we will show that this adjustment effectively removes the bias in our subgraph counting. We then use the value of $\tilde{a}_{i,c_{t}}$ to estimate the number of subgraphs associated with each node. The calculation method varies depending on the specific subgraph. In Line 4 of the algorithm, we detail the process for triangle counting. However, we can modify this line to calculate other graph statistics. Following this, in Line 5, we apply either the local Laplacian mechanism or the smooth sensitivity mechanism to obfuscate the result from Line 4 before sending it to the central server. At Line 6, the central server aggregates the results received from all users and publishes the summation as the final counting result.
It is crucial to understand that we do not assume the degree $d_{i}$ is public. Rather, we utilize an estimation of $d_{i}$, referred to as $\tilde{d}_{i}$, which is disclosed in the initial step of our mechanism, for the calculation of $\tilde{\sigma}_{i}$ at Line 2 of the algorithm.
1
2Function Counting
 Input: Graph $G=(V,E)$, grouped and obfuscated adjacency list $(a^{\prime}_{i,c_{1}},\dots,a^{\prime}_{i,c_{m}})$, privacy budget $\varepsilon_{2}$, estimated degree $\tilde{d_{i}}$ of each user.
 Output: The estimated number of triangles in the graph
3 [User $i$] Calculate $\tilde{\sigma}_{i}=\frac{s-1}{ms-s}\tilde{d_{i}}+\frac{1}{e^{\varepsilon^{\prime}}-1}\frac{ms-1}{m-1}$, $\omega_{i}=\frac{e^{\varepsilon^{\prime}}+1}{e^{\varepsilon^{\prime}}-1}\cdot\frac{ms-1}{m-1}$;
4 [User $i$] For $1\leq t\leq m$, calculate $\tilde{a}_{i,c_{t}}=\omega_{i}\cdot a^{\prime}_{i,c_{t}}-\tilde{\sigma}_{i}$ ;
5 [User $i$] Let $W_{i}:=\{(j,k):j<k<i\text{ and }\{v_{i},v_{j}\},\{v_{i},v_{k}\}\in E\}$. Calculate $|\tilde{\Delta}_{i}|=\sum\limits_{(j,k)\in W_{i}}\tilde{a}_{k,h_{k}(j)}$;
6 [User $i$] Obfuscate the value of $|\tilde{\Delta}_{i}|$ using the local Laclacian mechanism or the smooth sensitivity mechanism with privacy budget $\varepsilon_{2}$. Then, submit the result to the central server;
7 [Server] Publish the number of triangle as $\sum_{i}|\tilde{\Delta}_{i}|$
8
Algorithm 3 The third step of our method, counting step, when it is used for publishing the number of triangles
5 Theoretical Analysis of Our Mechanism
5.1 Privacy
We begin by analyzing the privacy guarantees of the group randomized response step (Step 2) in the following lemma. This analysis leverages the amplification by subsampling technique (Balle et al., 2018), defined as follows:
Lemma 1 (Amplification by Sub-Sampling under Edge Differential Privacy (Balle et al., 2018)).
Let $s>0$, and define $m=\lceil n/s\rceil$. Consider a randomized sampling function $\mathcal{S}:\{0,1\}^{n}\rightarrow\{0,1\}^{m}$ such that, given an input $(a_{1},\dots,a_{n})\in\{0,1\}^{n}$, the output is $\mathcal{S}(a_{1},\dots,a_{n})=(a^{\prime}_{1},\dots,a^{\prime}_{m})$, where each $a^{\prime}_{i}$ is determined as follows:
  • For each $1\leq i<m$, the output $a^{\prime}_{i}$ is sampled uniformly at random from the $i$-th block of size $s$, i.e., from indices $\{(i-1)s+1,\dots,is\}$. That is, for each $j\in\{(i-1)s+1,\dots,is\}$, we have $a^{\prime}_{i}=a_{j}$ with probability $1/s$.
  • For $i=m$, let the final block be $\{(m-1)s+1,\dots,n\}$, which may contain fewer than $s$ elements. Then, for each $j\in\{(m-1)s+1,\dots,n\}$, we have $a^{\prime}_{m}=a_{j}$ with probability $1/s$, and $a^{\prime}_{m}=0$ with probability $\frac{ms-n}{s}$.
[…]
$e^{-\varepsilon^{\prime}}\leq\frac{\Pr[\mathcal{M}(a)\in S]}{\Pr[\mathcal{M}(\hat{a})\in S]}\leq e^{\varepsilon^{\prime}}.$ (2)
[…]
$e^{-\varepsilon_{s}}\leq\frac{\Pr[\mathcal{M}(\mathcal{S}(a))\in S]}{\Pr[\mathcal{M}(\mathcal{S}(\hat{a}))\in S]}\leq e^{\varepsilon_{s}},$ (3)
where $\varepsilon_{s}$ is given by $\varepsilon_{s}=\ln\left(1+\frac{e^{\varepsilon^{\prime}}-1}{s}\right).$
The amplification by sub-sampling technique was originally proposed in the context of central differential privacy, where each $a_{i}$ represents sensitive information of user $i$. It has not been previously applied to either local differential privacy or graph statistics. The assumption in (2) requires that $\mathcal{M}$ satisfies $\varepsilon^{\prime}$-differential privacy, and the conclusion in (3) shows that the composed mechanism achieves $\varepsilon_{s}$-differential privacy due to the sampling. To the best of our knowledge, we are the first to apply amplification by sub-sampling in the setting of local differential privacy and graph statistics. We consider this application to be one of the main contributions of our work.
The privacy guarantee of the group randomized response step (Step 2) of our algorithm follows from Lemma 1.
Lemma 2.
The privacy budget of the group randomized response step is $\varepsilon_{1}$.
Proof.
We begin by observing that the computation of $a_{i,c_{1}},\dots,a_{i,c_{m}}]$ in Lines 4–6 of Algorithm 2 can be viewed as a sampling procedure $\mathcal{S}$, as described in Lemma 1, followed by a reordering step, which does not affect the privacy guarantees of the mechanism. The randomized response mechanism applied in Line 7 satisfies the privacy bound given in (2). Consequently, the overall procedure in Algorithm 2, which composes sampling with randomized response, satisfies equation (3), thereby ensuring $\varepsilon_{s}$-edge local differential privacy as defined in Definition 5. Substituting $\varepsilon^{\prime}=\ln(1+s(e^{\varepsilon_{1}}-1))$ into the expression for $\varepsilon_{s}=\ln(1+\frac{e^{\varepsilon^{\prime}}-1}{s})$, we conclude that $\varepsilon_{s}=\varepsilon_{1}$ and the privacy budget of the group randomized response step is $\varepsilon_{1}$. ∎
We provide the following example to illustrate how this amplification is achieved.
Example 1.
Consider the publication of node $\nu_{0}$’s adjacency list using the group randomized response. We will analyze two cases: (1) $\nu_{0}$ is connected to none of the n1n-1 other nodes, and (2) $\nu_{0}$ is connected only to $\nu_{1}$. Additionally, since the groups created by the hashing functions are public information and independent of the private adjacency list, we assume that $\nu_{1}$ is in $S_{0,t}$ for both cases.
Let $p_{1}$ and $p_{2}$ represent the probabilities that $a^{\prime}_{i,c_{t}}=1$ in scenarios (1) and (2), respectively, when using the randomized response mechanism with a privacy budget $\varepsilon^{\prime}$. In case (1), $a_{i,c_{t}}$ is never equal to 1, as no nodes are connected to $\nu_{0}$. By the property of randomized response, $p_{1}=\frac{1}{1+e^{\varepsilon^{\prime}}}$. In case (2), $a_{i,c_{t}}$ is equal to 1 only when $\nu_{1}$ is selected as the representative of group $S_{0,t}$ (note that this information is not disclosed). Thus, $a_{i,c_{t}}=1$ with probability $\frac{1}{s}$. We have $p_{2}=\frac{s-1}{s}\cdot\frac{1}{1+e^{\varepsilon^{\prime}}}+\frac{1}{s}\cdot\frac{e^{\varepsilon^{\prime}}}{1+e^{\varepsilon^{\prime}}}$. We then derive that $\frac{p_{2}}{p_{1}}=1+\frac{1}{s}\cdot\left(e^{\varepsilon^{\prime}}-1\right)$. By setting $\varepsilon_{1}$ such that $\varepsilon^{\prime}=\ln(1+s(e^{\varepsilon_{1}}-1))$, we obtain $\frac{p_{2}}{p_{1}}=e^{\varepsilon_{1}}$. This shows that we achieve $\varepsilon_{1}$-differential privacy, with $\varepsilon_{1}\approx\varepsilon^{\prime}/s$, demonstrating the privacy budget amplification.
The final result of our privacy analysis is shown in the following theorem:
Theorem 2.
The privacy budget of our mechanism is $\varepsilon_{0}+\varepsilon_{1}+\varepsilon_{2}$.
Proof.
We directly obtain this result from Lemma 2 and the composition theorem (Theorem 1). ∎
5.2 Expected Value of $\tilde{a}_{i,c_{t}}$
In the subsequent theorem, we discuss the expected value of $\tilde{a}_{i,c_{t}}$. The theorem reveals that when an estimate from the counting phase is a linear composition of $\tilde{a}_{i,c_{t}}$, it remains unbiased. The triangle count estimation in Algorithm 3 is an example of these estimations.
In the following, we will restrict our analysis to the values of $\tilde{a}_{j,h_{j}(k)}$ where $j>k$. This does not pose any limitation, as $a_{j,k}$ can always be estimated using either $\tilde{a}_{j,h_{j}(k)}$ or $\tilde{a}_{k,h_{k}(j)}$. Therefore, for the remainder of this section, we will assume that $j>k$.